# BotArtist: Generic approach for bot detection in Twitter via semi-automatic machine learning pipeline

### v4

@@ Abstract
Twitter, as one of the most popular social networks, provides a platform for communication and online discourse. Unfortunately, it has also become a target for bots and fake accounts, resulting in the spread of false information and manipulation. This paper introduces a semi-automatic machine learning pipeline (SAMLP) designed to address the challenges correlated with machine learning model development. Through this pipeline, we develop a comprehensive bot detection model named BotArtist, based on user profile features. SAMLP leverages nine distinct publicly available datasets to train the BotArtist model. To assess BotArtist’s performance against current state-of-the-art solutions, we select 35 existing Twitter bot detection methods, each utilizing a diverse range of features. Our comparative evaluation of BotArtist and these existing methods, conducted across nine public datasets under standardized conditions, reveals that the proposed model outperforms existing solutions by almost 10%, in terms of F1-score, achieving an average score of 83.19 and 68.5 over specific and general approaches respectively. As a result of this research, we provide a dataset of the extracted features combined with BotArtist predictions over the 10.929.533 Twitter user profiles, collected via Twitter API during the 2022 Russo-Ukrainian War, over a 16-month period. This dataset was created in collaboration with (Shevtsov et al., 2022a) where the original authors share anonymized tweets on the discussion of the Russo-Ukrainian war with a total amount of 127.275.386 tweets. The combination of the existing text dataset and the provided labeled bot and human profiles will allow for the future development of a more advanced bot detection large language model in the post-Twitter API era.
1 Introduction
@@
Twitter, one of the most popular social networks, with millions of active users, is used for news dissemination, political discussions, and social interactions. However, the platform has also been plagued by bots and fake accounts that are used to manipulate and spread false information. According to the research community, the usage of manipulation techniques implemented with the use of bot accounts is registered during diverse popular topic discussions. More specifically, studies show that bot accounts are involved in diverse political discussions, for instance, around 2016 in countries like the US, Germany, Sweden, France, Spain, etc. (Golovchenko et al., 2020; Badawy et al., 2018; Howard et al., 2016; Shevtsov et al., 2022b, 2023; Neudert et al., 2017; Pastor-Galindo et al., 2020; Bradshaw et al., 2017; Fernquist et al., 2018; Castillo et al., 2019; Rossi et al., 2020). Furthermore, propagandizing and false news is identified during the vaccination debate (Broniatowski et al., 2018) and more recent examples of the COVID-19 pandemic (Shahi et al., 2021; Ferrara, 2020; Yang et al., 2021). This high activity of bot accounts raises concerns in the research community and online social media platforms about the integrity of the shared information.
@@
Currently, the research community offers a variety of ML and NN-based bot detection methods. Each of these provided methods excels at addressing specific bot detection scenarios, yielding optimal solutions for particular use cases. Unfortunately, the performance of these provided methods significantly diminishes in more general bot detection scenarios, encompassing different periods, discussion topics, and languages, among others. Recent studies have shown that existing methods still do not achieve flawless Twitter bot detection, regardless of the number of user characteristics that they incorporate (Feng et al., 2022a). In addition, a large portion of existing ML bot detection methods ignore well-known optimization procedures (including feature selection and dynamic hyper-parameter fine-tuning) providing significant space for improvement.
@@
In our research, we address this challenge by developing a semi-automatic machine learning pipeline (SAMLP111GitHub repository: TBD ) for constructing a generic bot detection model (BotArtist222GitHub repository: TBD ). The presented pipeline resolves popular issues during the creation of the classification model, including recursive hyperparameter fine-tuning during the feature selection phase. This approach guarantees the reduction of non-relevant features which sequentially reduces the noise of initial data. Additionally, our pipeline considers the class imbalance during feature selection and model fine-tuning, making the model suitable for real-case applications. Next, we compare this developed model with 35 currently available state-of-the-art approaches using nine publicly available datasets. This research yields two distinct evaluation scenarios: one specific to each dataset, where each model is trained and tested separately on each of them, and a more generic approach, where the models are trained and tested on the combined dataset comprising all nine unique datasets.
@@
Except from the source code of SAMLP and BotArtist, we make available the model outcome usable for research. For this purpose, we collect a large volume of users correlated to the already existing textual shared dataset and add an additional layer to textual information via user prediction. We then release 10.929.533 user profiles with BotArtist predictions 333Zenodo repository: TBD , correlated to 127.275.386 tweets (Shevtsov et al., 2022a).
2 Related Work
@@ 2.1 Feature based
Methods falling into this category apply feature engineering based on information extracted from Twitter user profiles and their activity patterns. Researchers use traditional machine learning or neural network classification algorithms for bot detection. These approaches employ different sets of user characteristics as feature sets, including user metadata (Kudugunta and Ferrara, 2018), tweets (Miller et al., 2014), user name (Beskow and Carley, 2019), description (Hayawi et al., 2022), temporal patterns (Mazza et al., 2019), and follow relationships (Feng et al., 2021a). Moreover, some of them aim to enhance the scalability of feature-based methods (Yang et al., 2020; Abreu et al., 2020), discover unknown bot accounts using correlation-based techniques (Chavoshi et al., 2016), and improve the trade-off between precision and recall in bot detection (Morstatter et al., 2016). However, the creators of bot accounts are increasingly aware of the features utilized by the research community, leading to novel bot implementations designed to evade detection based on known features (Cresci, 2020). Consequently, existing feature-based methods face several challenges in the accurate detection of these new bot accounts (Feng et al., 2021a).
2.2 Text based
@@
Text-based methods rely on natural language processing (NLP) techniques for Twitter bot detection, extracting characteristics from posted content (tweets) and user profile descriptions. Approaches in this category include sequence fingerprint (Cresci et al., 2016), word embeddings (Wei and Nguyen, 2019), recurrent neural networks (RNN) (Kudugunta and Ferrara, 2018), attention mechanisms (Feng et al., 2021a), transformers (Guo et al., 2021a; Liu et al., 2019; Raffel et al., 2020) and the adoption of pre-trained language models for encoding tweets (Dukić et al., 2020). Researchers have also combined tweet representations with user profile characteristics (Cai et al., 2017; Efthimion et al., 2018; Kantepe and Ganiz, 2017; Miller et al., 2014; Varol et al., 2017; Kouvela et al., 2020; Lee et al., 2011; Echeverria et al., 2018), employed unsupervised machine learning models (Feng et al., 2021a), and addressed multilingual context issues (Knauth, 2019). Despite extensive existing research studies and impressive performance in text-based approaches, new bot accounts can still evade detection by sharing stolen content from genuine users (Cresci, 2020). Additionally, recent work has demonstrated that relying solely on textual information is insufficient for robust and accurate bot detection (Feng et al., 2021b).
2.3 Graph based
@@
The graph-based category of bot detection methods combines geometric deep neural networks with graph analytics. Current implementations leverage techniques such as node centrality (Dehghan et al., 2023), node representation learning (Pham et al., 2022), graph neural networks (GNNs) (Ali Alhosseini et al., 2019; Moghaddam and Abbaspour, 2022), and heterogeneous GNNs (Feng et al., 2021b) to conduct graph-based bot detection. In recent research, the authors have borrowed ideas from other categories to merge different approaches, such as combining multiple methods (Guo et al., 2021b; Yang et al., 2022; Knauth, 2019; Beskow and Carley, 2020; Rodríguez-Ruiz et al., 2020; Magelinski et al., 2020; Yang et al., 2013; Kipf and Welling, 2016, 2016; Veličković et al., 2017; Lv et al., 2021). Furthermore, novel GNN architectures have been proposed to exploit heterogeneity in the Twitter network (Feng et al., 2022b). Generally, these approaches hold significant promise for Twitter bot detection.
@@
Unfortunately, existing Twitter bot detection methods have their limitations. Simplified feature-based approaches struggle to generalize and provide robust results, while more complex methods based on text or graph characteristics require intensive computational resources and large datasets for model development. Furthermore, as shown in (Shevtsov et al., 2024) the complex embedding-based model tends to perfectly capture the training data patterns, but lacks generalizability and requires frequent retraining over multiple datasets to keep high performance. These limitations, coupled with the recent announcement of the Twitter API monetization (Twitter, 2023), make many existing methods expensive to maintain or nearly impossible to operate daily. In this study, we address these challenges and present robust and accurate Twitter bot detection methods, based on a lightweight set of features. Our methods require only a single user profile object (Twitter API v1.1 or v2) to provide user prediction without additional Twitter API requests, thereby significantly reducing operational costs.
3 Datasets
@@
For this research paper, we collect nine well-known publicly available datasets (Feng et al., 2021c). All selected datasets already contain ground truth labels, primarily obtained through manual analysis or crowd-sourcing. For simplicity, we label the selected datasets as follows: C-15 (Cresci et al., 2015), G-17 (Gilani et al., 2017), C-17 (Cresci et al., 2017a, b), M-18 (Yang et al., 2020), C-S-18 (Cresci et al., 2018, 2019), C-R-19 (Mazza et al., 2019), B-F-19 (Yang et al., 2019), TwiBot-20 (Feng et al., 2021c), and TwiBot-22 (Feng et al., 2022a). In Table 1, we present the information provided in each dataset, along with the volume of normal and bot accounts.
@@
Furthermore, in collaboration with (Shevtsov et al., 2022a) collect 10.929.533 Twitter profiles correlated with 127.275.386 publicly available tweets related to the public discussion topic of the 2022 Russo-Ukrainian War. Our collection of user profiles is based on the monitoring of selected topics starting from February 23, 2022, till June 23, 2023. The shared datasets contain a set of extracted features in an anonymized form of a CSV file and contain only preprocessed numerical features to protect user information. The provided dataset also provides anonymized user IDs which are identically correlated with publicly available user tweet dataset (Shevtsov et al., 2022a).
4 Methodology
@@ 4.1 Feature Selection
4.2 Model Fine-Tuning
Having reduced the problem’s dimensionality through feature selection, we now focus on identifying the model and its configuration (hyper-parameter set) that can provide accurate predictions over unseen data samples. We select three well-known machine learning classification models: SVM (Boser et al., 1992), RandomForest (Breiman, 2001), and state-of-the-art XGBoost (Extreme Gradient Boosting) (Chen et al., 2015).
@@ 4.2 Model Fine-Tuning
Afterward, we select the best classifier and configuration based on the average validation (F1-score). This model configuration becomes the final model, which will be evaluated over the hold-out (testing) data portion. In the final stage of the model development, we provide model explainability using the SHAP game-theoretical approach (Shapley, 1953). This allows us to describe the reasons behind the model’s decisions and highlight the distinctions between classes.
@@
4.3 Feature Extraction
As mentioned earlier, in this study we pursue the creation of a simple model that in comparison with other bot detection systems- does not require a large volume of information or features for accurate prediction. In our case, the number of Twitter profiles triggers our attention. The majority of social media bot accounts are created for content promotion and increasing the attention of other registered users. To maximize this goal, bot accounts aim for the expansion of social audience. Social expansion is very difficult to achieve without high activity, such as tweets, re-tweets, comments, etc. In comparison with profile-based features, the extraction and utilization of textual or graph features can result in instability, causing a significant performance drop across various periods or discussion topic datasets (Shevtsov et al., 2024). Consequently, we need to extract as much information as possible from the user profiles to identify extremely active accounts. The difference between very popular social media accounts (also known as celebrities), and bot accounts is that they try to post and share as much content as possible during the day. Due to the differences between Twitter API v1.1 and v2, we have devised a method for constructing features in a way that allows the extraction of selected information from both v1.1 and v2 without any additional manipulation. The profile features we have extracted are categorized into three distinct groups: count, boolean, and real-valued.
@@ 4.3 Feature Extraction
5 Experimental Results
According to the methodology of SAMLP described earlier, we successfully develop a semi-automated machine learning pipeline capable of constructing classification models (both binary and multiclass) for both balanced and imbalanced datasets. Using this pipeline, we create and test our custom implementation called BotArtist for Twitter bot detection, utilizing nine well-known datasets. To evaluate BotArtist’s performance, in comparison to 35 other existing research approaches, we employ the most comprehensive Twitter bot detection benchmark available (Feng et al., 2022a), where various methods are compared using identical training/validation and testing data portions. This benchmarking allows us to make a fair comparison between our approach and established bot detection methods.
@@ 5 Experimental Results
5.1 Models comparison
@@
To measure the performance of the trained methods, we evaluate them by testing on portions of each dataset and also on the union of all the dataset testing portions. These approaches allow us to not only identify the general performance of each bot detection method but also gain insights into performance variations across each specific dataset. To achieve this, we selectively choose the most accurate models from Table 3, in addition to the Botometer model (Yang et al., 2022). The Botometer model is included because it has already been trained on a diverse range of datasets and is expected to perform well in general-case scenarios.

### v5

@@ Abstract
Twitter, as one of the most popular social networks, provides a platform for communication and online discourse. Unfortunately, it has also become a target for bots and fake accounts, resulting in the spread of false information and manipulation. This paper introduces a semi-automatic machine learning pipeline (SAMLP) designed to address the challenges associated with machine learning model development. Through this pipeline, we develop a comprehensive bot detection model named BotArtist, based on user profile features.
SAMLP leverages nine distinct publicly available datasets to train the BotArtist model. To assess BotArtist’s performance against current state-of-the-art solutions, we evaluate 35 existing Twitter bot detection methods, each utilizing a diverse range of features. Our comparative evaluation of BotArtist and these existing methods, conducted across nine public datasets under standardized conditions, reveals that the proposed model outperforms existing solutions by almost 10% in terms of F1-score, achieving an average score of 83.19% and 68.5% over specific and general approaches, respectively.
As a result of this research, we provide one of the largest labeled Twitter bot datasets. The dataset contains extracted features combined with BotArtist predictions for 10,929,533 Twitter user profiles, collected via Twitter API during the 2022 Russo-Ukrainian War over a 16-month period. This dataset was created based on (Shevtsov et al. 2022b) where the original authors share anonymized tweets discussing the Russo-Ukrainian war, totaling 127,275,386 tweets. The combination of the existing textual dataset and the provided labeled bot and human profiles will enable future development of more advanced bot detection large language models in the post-Twitter API era.
Introduction
@@
Twitter, one of the most popular social networks, with millions of active users, is used for news dissemination, political discussions, and social interactions. However, the platform has also been plagued by bots and fake accounts that are used to manipulate and spread false information. According to the research community, the usage of manipulation techniques implemented with the use of bot accounts is registered during diverse popular topic discussions. More specifically, studies show that bot accounts are involved in diverse political discussions, for instance, around 2016 in countries like the US, Germany, Sweden, France, Spain, etc. (Golovchenko et al. 2020; Badawy, Ferrara, and Lerman 2018; Howard, Kollanyi, and Woolley 2016; Shevtsov et al. 2022a, 2023; Neudert, Kollanyi, and Howard 2017; Pastor-Galindo et al. 2020; Bradshaw et al. 2017; Fernquist, Kaati, and Schroeder 2018; Castillo et al. 2019; Rossi et al. 2020). Furthermore, propagandizing and false news is identified during the vaccination debate (Broniatowski et al. 2018) and more recent examples of the COVID-19 pandemic (Shahi, Dirkson, and Majchrzak 2021; Ferrara 2020; Yang et al. 2021). This high activity of bot accounts raises concerns in the research community and online social media platforms about the integrity of the shared information.
@@
Currently, the research community offers a variety of ML and NN-based bot detection methods. Each of these provided methods excels at addressing specific bot detection scenarios, yielding optimal solutions for particular use cases. Unfortunately, the performance of these provided methods significantly diminishes in more general bot detection scenarios, encompassing different periods, discussion topics, and languages, among others. Recent studies have shown that existing methods still do not achieve flawless Twitter bot detection, regardless of the number of user characteristics that they incorporate (Feng et al. 2022b). In addition, a large portion of existing ML bot detection methods ignore well-known optimization procedures (including feature selection and dynamic hyper-parameter fine-tuning) providing significant space for improvement.
@@
In our research, we address this challenge by developing a semi-automatic machine learning pipeline (SAMLP111GitHub repository: https://github.com/alexdrk14/SAMLP ) for constructing a generic bot detection model (BotArtist222GitHub repository: https://github.com/alexdrk14/BotArtist). The presented pipeline resolves popular issues during the creation of the classification model, including recursive hyperparameter fine-tuning during the feature selection phase. This approach guarantees the reduction of non-relevant features which sequentially reduces the noise of initial data. Additionally, our pipeline considers the class imbalance during feature selection and model fine-tuning, making the model suitable for real-case applications. Next, we compare this developed model with 35 currently available state-of-the-art approaches using nine publicly available datasets. This research yields two distinct evaluation scenarios: one specific to each dataset, where each model is trained and tested separately on each of them, and a more generic approach, where the models are trained and tested on the combined dataset comprising all nine unique datasets.
@@
Except from the source code of SAMLP and BotArtist, we make available the model outcome usable for research. For this purpose, we collect a large volume of users correlated to the already existing textual shared dataset and add an additional layer to textual information via user prediction. We then release 10.929.533 user profiles with BotArtist predictions 333Zenodo repository: https://zenodo.org/records/11203900, correlated to 127.275.386 tweets (Shevtsov et al. 2022b).
| Dataset | C-15 | G-17 | C-17 | M-18 | C-S-18 | C-R-19 | B-F-19 | TwiBot-20 | TwiBot-22 |
| # Total User | 5,301 | 2,484 | 14,368 | 50,538 | 13,276 | 693 | 518 | 229,580 | 1,000,000 |
| # Human | 1,950 | 1,394 | 3,474 | 8,092 | 6,174 | 340 | 380 | 5,237 | 860,057 |
| # Bot | 3,351 | 1,090 | 10,894 | 42,446 | 7,102 | 353 | 138 | 6,589 | 139,943 |
| # Total Tweet | 2,827,757 | 0 | 6,637,615 | 0 | 0 | 0 | 0 | 33,488,192 | 88,217,457 |
| # Human Tweet | 2,631,730 | 0 | 2,839,361 | 0 | 0 | 0 | 0 | 927,292 | 81,250,102 |
| # Bot Tweet | 196,027 | 0 | 3,798,254 | 0 | 0 | 0 | 0 | 1,072,496 | 6,967,355 |
| # Graph Edges | 7,086,134 | 0 | 6,637,615 | 0 | 0 | 0 | 0 | 33,488,192 | 170,185,937 |
Related Work
@@
Methods falling into this category apply feature engineering based on information extracted from Twitter user profiles and their activity patterns. Researchers use traditional machine learning or neural network classification algorithms for bot detection. These approaches employ different sets of user characteristics as feature sets, including user metadata (Kudugunta and Ferrara 2018), tweets (Miller et al. 2014), user name (Beskow and Carley 2019), description (Hayawi et al. 2022), temporal patterns (Mazza et al. 2019), and follow relationships (Feng et al. 2021a). Moreover, some of them aim to enhance the scalability of feature-based methods (Yang et al. 2020; Abreu, Ralha, and Gondim 2020), discover unknown bot accounts using correlation-based techniques (Chavoshi, Hamooni, and Mueen 2016), and improve the trade-off between precision and recall in bot detection (Morstatter et al. 2016). However, the creators of bot accounts are increasingly aware of the features utilized by the research community, leading to novel bot implementations designed to evade detection based on known features (Cresci 2020). Consequently, existing feature-based methods face several challenges in the accurate detection of these new bot accounts (Feng et al. 2021a).
Text based
@@
Text-based methods rely on natural language processing (NLP) techniques for Twitter bot detection, extracting characteristics from posted content (tweets) and user profile descriptions. Approaches in this category include sequence fingerprint (Cresci et al. 2016), word embeddings (Wei and Nguyen 2019), recurrent neural networks (RNN) (Kudugunta and Ferrara 2018), attention mechanisms (Feng et al. 2021a), transformers (Guo et al. 2021a; Liu et al. 2019; Raffel et al. 2020) and the adoption of pre-trained language models for encoding tweets (Dukić, Keča, and Stipić 2020). Researchers have also combined tweet representations with user profile characteristics (Cai, Li, and Zengi 2017; Efthimion, Payne, and Proferes 2018; Kantepe and Ganiz 2017; Miller et al. 2014; Varol et al. 2017; Kouvela, Dimitriadis, and Vakali 2020; Lee, Eoff, and Caverlee 2011; Echeverria et al. 2018), employed unsupervised machine learning models (Feng et al. 2021a), and addressed multilingual context issues (Knauth 2019). Despite extensive existing research studies and impressive performance in text-based approaches, new bot accounts can still evade detection by sharing stolen content from genuine users (Cresci 2020). Additionally, recent work has demonstrated that relying solely on textual information is insufficient for robust and accurate bot detection (Feng et al. 2021c).
Graph based
@@
The graph-based category of bot detection methods combines geometric deep neural networks with graph analytics. Current implementations leverage techniques such as node centrality (Dehghan et al. 2023), node representation learning (Pham et al. 2022), graph neural networks (GNNs) (Ali Alhosseini et al. 2019; Moghaddam and Abbaspour 2022), and heterogeneous GNNs (Feng et al. 2021c) to conduct graph-based bot detection. In recent research, the authors have borrowed ideas from other categories to merge different approaches, such as combining multiple methods (Guo et al. 2021b; Yang, Ferrara, and Menczer 2022; Knauth 2019; Beskow and Carley 2020; Rodríguez-Ruiz et al. 2020; Magelinski, Beskow, and Carley 2020; Yang, Harkreader, and Gu 2013; Kipf and Welling 2016, 2016; Veličković et al. 2017; Lv et al. 2021). Furthermore, novel GNN architectures have been proposed to exploit heterogeneity in the Twitter network (Feng et al. 2022a). Generally, these approaches hold significant promise for Twitter bot detection.
@@
Unfortunately, existing Twitter bot detection methods have their limitations. Simplified feature-based approaches struggle to generalize and provide robust results, while more complex methods based on text or graph characteristics require intensive computational resources and large datasets for model development. Furthermore, as shown in (Shevtsov et al. 2024) the complex embedding-based model tends to perfectly capture the training data patterns, but lacks generalizability and requires frequent retraining over multiple datasets to keep high performance. These limitations, coupled with the recent announcement of the Twitter API monetization (Twitter 2023), make many existing methods expensive to maintain or nearly impossible to operate daily. In this study, we address these challenges and present robust and accurate Twitter bot detection methods, based on a lightweight set of features. Our methods require only a single user profile object (Twitter API v1.1 or v2) to provide user prediction without additional Twitter API requests, thereby significantly reducing operational costs.
Datasets
@@
For this research paper, we collect nine well-known publicly available datasets (Feng et al. 2021b). All selected datasets already contain ground truth labels, primarily obtained through manual analysis or crowd-sourcing. For simplicity, we label the selected datasets as follows: C-15 (Cresci et al. 2015), G-17 (Gilani et al. 2017), C-17 (Cresci et al. 2017b, a), M-18 (Yang et al. 2020), C-S-18 (Cresci et al. 2018, 2019), C-R-19 (Mazza et al. 2019), B-F-19 (Yang et al. 2019), TwiBot-20 (Feng et al. 2021b), and TwiBot-22 (Feng et al. 2022b). In Table 1, we present the information provided in each dataset, along with the volume of normal and bot accounts.
@@
Furthermore, in collaboration with (Shevtsov et al. 2022b) collect 10.929.533 Twitter profiles correlated with 127.275.386 publicly available tweets related to the public discussion topic of the 2022 Russo-Ukrainian War. Our collection of user profiles is based on the monitoring of selected topics starting from February 23, 2022, till June 23, 2023. The shared datasets contain a set of extracted features in an anonymized form of a CSV file and contain only preprocessed numerical features to protect user information. The provided dataset also provides anonymized user IDs which are identically correlated with publicly available user tweet dataset (Shevtsov et al. 2022b).
Methodology
@@
Model Fine-Tuning
Having reduced the problem’s dimensionality through feature selection, we now focus on identifying the model and its configuration (hyper-parameter set) that can provide accurate predictions over unseen data samples. We select three well-known machine learning classification models: SVM (Boser, Guyon, and Vapnik 1992), RandomForest (Breiman 2001), and state-of-the-art XGBoost (Extreme Gradient Boosting) (Chen et al. 2015).
@@
Afterward, we select the best classifier and configuration based on the average validation (F1-score). This model configuration becomes the final model, which will be evaluated over the hold-out (testing) data portion. In the final stage of the model development, we provide model explainability using the SHAP game-theoretical approach (Shapley 1953). This allows us to describe the reasons behind the model’s decisions and highlight the distinctions between classes.
@@
Described experiments were conducted in the single machine with AMD Ryzen 9 CPU (16 cores/32 threads) with a total of 64 GB DDR4 system memory. For the graphical unit based computations the single unit of Nvidia RTX 3080 where used with total memory of 12 GB.
| Feature | Type | Feature | Type | Calculation |
| name_len | count | screen_name_sim | real-valued | Jaccard of (screen_name, user_name) |
| screen_name_len | count | foll_friends | real-valued | follower / friends |
| description_len | count | age | real-valued | created_at / collection date |
| listed | count | listed_by_age | real-valued | listed / account age |
| statuses | count | statuses_by_age | real-valued | statuses / account age |
| followers | count | followers_by_age | real-valued | followers / account age |
| following | count | following_by_age | real-valued | friends / account age |
| name_upper_len | count | name_upper_pcnt | real-valued | percentage of upper case |
| name_lower_len | count | name_lower_pcnt | real-valued | percentage of lower case |
| name_digits_len | count | name_digits_pcnt | real-valued | percentage of digits |
| name_special_len | count | name_special_pcnt | real-valued | percentage of other characters |
| screen_name_upper_len | count | screen_name_upper_pcnt | real-valued | percentage of upper case |
| screen_name_lower_len | count | screen_name_lower_pcnt | real-valued | percentage of lower case |
| screen_name_digits_len | count | screen_name_digits_pcnt | real-valued | percentage of digits |
| screen_name_special_len | count | screen_name_special_pcnt | real-valued | percentage of other characters |
| description_upper_len | count | description_upper_pcnt | real-valued | percentage of upper case |
| description_lower_len | count | description_lower_pcnt | real-valued | percentage of lower case |
| description_digits_len | count | description_digits_pcnt | real-valued | percentage of digits |
| description_special_len | count | description_special_pcnt | real-valued | percentage of other characters |
| description_urls | count | name_entropy | real-valued | entropy of user name |
| description_mentions | count | screen_name_entropy | real-valued | entropy of the screen name |
| description_hashtags | count | has_location | boolean | |
| total_urls | count | has_profile_image | boolean | |
| protected | boolean | has_profile_url | boolean | |
| verified | boolean |
Feature Extraction
As mentioned earlier, in this study we pursue the creation of a simple model that in comparison with other bot detection systems- does not require a large volume of information or features for accurate prediction. In our case, the number of Twitter profiles triggers our attention. The majority of social media bot accounts are created for content promotion and increasing the attention of other registered users. To maximize this goal, bot accounts aim for the expansion of social audience. Social expansion is very difficult to achieve without high activity, such as tweets, re-tweets, comments, etc. In comparison with profile-based features, the extraction and utilization of textual or graph features can result in instability, causing a significant performance drop across various periods or discussion topic datasets (Shevtsov et al. 2024). Consequently, we need to extract as much information as possible from the user profiles to identify extremely active accounts. The difference between very popular social media accounts (also known as celebrities), and bot accounts is that they try to post and share as much content as possible during the day. Due to the differences between Twitter API v1.1 and v2, we have devised a method for constructing features in a way that allows the extraction of selected information from both v1.1 and v2 without any additional manipulation. The profile features we have extracted are categorized into three distinct groups: count, boolean, and real-valued.
@@
| Method | Type | C-15 | G-17 | C-17 | M-18 | C-S-18 | C-R-19 | B-F-19 | TB-20 | TB-22 | Average |
| SGBot | F | 77.9 | 72.1 | 94.6 | 99.5 | 82.3 | 82.7 | 49.6 | 84.9 | 36.6 | 75.57 |
| Kudugunta et al. | F | 75.3 | 49.8 | 91.7 | 94.5 | 50.9 | 49.2 | 49.6 | 47.3 | 51.7 | 62.22 |
| Hayawi et al. | F | 85.6 | 34.7 | 93.8 | 91.5 | 60.8 | 60.9 | 20.5 | 77.1 | 24.7 | 61.06 |
| BotHunter | F | 97.2 | 69.2 | 91.6 | 99.6 | 82.2 | 82.9 | 49.6 | 79.1 | 23.5 | 74.98 |
| NameBot | F | 83.4 | 44.8 | 85.7 | 91.6 | 61.1 | 67.5 | 38.5 | 65.1 | 0.5 | 59.80 |
| Abreu et al. | F | 76.4 | 66.7 | 95.0 | 97.9 | 76.9 | 83.5 | 53.8 | 77.1 | 53.4 | 75.63 |
| BotArtist | F | 98.3 | 76.1 | 97.0 | 99.7 | 80.6 | 88.3 | 68.4 | 82.2 | 58.2 | 83.19 |
| Cresci | T | 1.17 | - | 22.8 | - | - | - | - | 13.7 | - | - |
| Wei | T | 82.7 | - | 78.4 | - | - | - | - | 57.3 | 53.6 | - |
| BGSRD | T | 90.8 | 35.7 | 86.3 | 90.5 | 58.2 | 41.1 | 13.0 | 70.0 | 21.1 | 56.30 |
| RoBERTa | T | 95.8 | - | 94.3 | - | - | - | - | 73.1 | 20.5 | - |
| T5 | T | 89.3 | - | 92.3 | - | - | - | - | 70.5 | 20.2 | - |
| Efthimion | FT | 94.1 | 5.2 | 91.8 | 95.9 | 68.2 | 71.7 | 0.0 | 67.2 | 27.5 | 57.95 |
| Kantepe | FT | 78.2 | - | 79.4 | - | - | - | - | 62.2 | 58.7 | - |
| Miller | FT | 83.8 | 59.9 | 86.8 | 91.1 | 56.8 | 43.6 | 0.0 | 74.8 | 45.3 | 60.23 |
| Varol | FT | 94.7 | - | - | - | - | - | - | 81.1 | 27.5 | - |
| Kouvela | FT | 98.2 | 66.6 | 99.1 | 98.2 | 80.4 | 81.1 | 28.1 | 86.5 | 30.0 | 74.24 |
| Santos | FT | 78.8 | 14.5 | 83.0 | 92.4 | 65.2 | 75.7 | 21.0 | 60.3 | - | - |
| Lee | FT | 98.6 | 67.8 | 99.3 | 97.9 | 82.5 | 82.7 | 50.3 | 80.0 | 30.4 | 76.61 |
| LOBO | FT | 98.8 | - | 97.7 | - | - | - | - | 80.8 | 38.6 | - |
| Moghaddam | FG | 73.9 | - | - | - | - | - | - | 79.9 | 32.1 | - |
| Alhosseini | FG | 92.2 | - | - | - | - | - | - | 72.0 | 38.1 | - |
| Knauth | FTG | 91.2 | 39.1 | 93.4 | 91.3 | 94.0 | 54.2 | 41.3 | 85.2 | 37.1 | 69.64 |
| FriendBot | FTG | 97.6 | - | 87.4 | - | - | - | - | 80.0 | - | - |
| SATAR | FTG | 95.0 | - | - | - | - | - | - | 86.1 | - | - |
| Botometer | FTG | 66.9 | 77.4 | 96.1 | 46.0 | 79.6 | 79.0 | 30.8 | 53.1 | 42.8 | 63.5 |
| Rodrifuez-Ruiz | FTG | 87.7 | - | 85.7 | - | - | - | - | 63.1 | 56.6 | - |
| GraphHist | FTG | 84.5 | - | - | - | - | - | - | 67.6 | - | - |
| EvolveBot | FTG | 90.1 | - | - | - | - | - | - | 69.7 | 14.1 | - |
| Dehghan | FTG | 88.3 | - | - | - | - | - | - | 76.2 | - | - |
| GCN | FTG | 97.2 | - | - | - | - | - | - | 80.8 | 54.9 | - |
| GAT | FTG | 97.6 | - | - | - | - | - | - | 85.2 | 55.8 | - |
| HGT | FTG | 96.9 | - | - | - | - | - | - | 88.2 | 39.6 | - |
| SimpleHGN | FTG | 97.5 | - | - | - | - | - | - | 88.2 | 45.4 | - |
| BotRGCN | FTG | 97.3 | - | - | - | - | - | - | 87.3 | 57.5 | - |
| RGT | FTG | 97.8 | - | - | - | - | - | - | 88.0 | 42.9 | - |
Experimental Results
According to the methodology of SAMLP described earlier, we successfully develop a semi-automated machine learning pipeline capable of constructing classification models (both binary and multiclass) for both balanced and imbalanced datasets. Using this pipeline, we create and test our custom implementation called BotArtist for Twitter bot detection, utilizing nine well-known datasets. To evaluate BotArtist’s performance, in comparison to 35 other existing research approaches, we employ the most comprehensive Twitter bot detection benchmark available (Feng et al. 2022b), where various methods are compared using identical training/validation and testing data portions. This benchmarking allows us to make a fair comparison between our approach and established bot detection methods.
@@
| Method | C-15 | G-17 | C-17 | M-18 | C-S-18 | C-R-19 | B-F-19 | TB-20 | TB-22 | Total | Average |
| BotArtist | 82.7 | 39.9 | 87.3 | 99.0 | 80.6 | 73.8 | 16.6 | 80.3 | 56.9 | 63.7 | 68.5 |
| Lee | 82.3 | 0.0 | 83.6 | 97.7 | 78.2 | 67.7 | 20.0 | 8.5 | 42.4 | 52.9 | 53.3 |
| Abreu | 84.4 | 0.3 | 80.1 | 88.4 | 67.1 | 40.8 | 11.7 | 15.6 | 29.0 | 40.4 | 46.3 |
| SGBot | 75.0 | 3.6 | 79.8 | 99.2 | 76.7 | 68.9 | 0.0 | 15.2 | 43.3 | 53.8 | 51.5 |
| BotHunter | 73.4 | 7.1 | 76.0 | 99.2 | 76.0 | 44.8 | 11.1 | 14.7 | 28.0 | 43.1 | 47.8 |
| Kouvela | 95.5 | 20.4 | 94.7 | 98.1 | 78.4 | 71.4 | 21.0 | 28.5 | 36.0 | 52.0 | 60.5 |
| Botometer | 66.9 | 77.4 | 96.1 | 46.0 | 79.6 | 79.0 | 30.8 | 53.1 | 42.8 | 45.3 | 63.5 |
Models comparison
@@
To measure the performance of the trained methods, we evaluate them by testing on portions of each dataset and also on the union of all the dataset testing portions. These approaches allow us to not only identify the general performance of each bot detection method but also gain insights into performance variations across each specific dataset. To achieve this, we selectively choose the most accurate models from Table 3, in addition to the Botometer model (Yang, Ferrara, and Menczer 2022). The Botometer model is included because it has already been trained on a diverse range of datasets and is expected to perform well in general-case scenarios.